OpenStudy (anonymous):

If (x+2)/2 = y/5, then which of the following must be true? x/2 = (y-2)/5 x/2 = (y-5)/5 (x+2)/5 = y/25

OpenStudy (phi):

you could use algebra to see if you can re-arrange the equation to get one of the choices. you could also test (x,y) pairs and see if any of the choices works. For example, if you choose x=0 and use that in (x+2)/2= y/5 (0+2)/2 = y/5 2/2 = y/5 or 1= y/5 y must be 5 if x is 0. See if (0,5) only works in one of the choices.

OpenStudy (anonymous):

Ok Thank You

OpenStudy (phi):

Did you get an answer ?

OpenStudy (anonymous):

Not Yet

OpenStudy (anonymous):

Can you possibly go into further explanation?

OpenStudy (phi):

in the first choice, replace x with 0 and y with 5 write down what you get

OpenStudy (anonymous):

Is it the last choice?

OpenStudy (anonymous):

This is soo confusing...